A penny is dropped into a tank of water's surface. It falls to the bottom according to the relation d=3.5t^2+35,where d is the depth of the water measured in meters and t is the time after the penny was dropped, measured in seconds. Complete the table of values for the relation d =3.5t^2 +35


1
Expert's answer
2020-11-01T14:56:16-0500

If t = 0 sec, then d=3.5 x 02 + 35 = 35 m

If t = 1 sec, then d=3.5 x 12 + 35 = 38.5 m

If t = 2 sec, then d=3.5 x 22 + 35 = 49 m

If t = 3 sec, then d=3.5 x 32 + 35 = 66.5 m

If t = 4 sec, then d=3.5 x 42 + 35 = 91 m

If t = 5 sec, then d=3.5 x 52 + 35 = 122.5 m

If t = 6 sec, then d=3.5 x 62 + 35 = 161 m

If t = 7 sec, then d=3.5 x 72 + 35 = 206.5 m

If t = 8 sec, then d=3.5 x 82 + 35 = 259 m

If t = 9 sec, then d=3.5 x 92 + 35 = 318.5 m

If t = 10 sec, then d=3.5 x 102 + 35 = 385 m